# Scintillation Properties of PSR B1133+16 Measured with Very Long Baseline Interferometry

### v1

@@ 5.2 Association with the Local Bubble
Figure 5: The differential dust extinction as a function of distance on the line of sight towards PSR B1133+16, based on the dust map of Edenhofer et al. (2024). The distances of the screens associated with the outer arc, middle arc, and inner arc are shown with green solid, orange dashed, and blue dashdashdotted lines with shaded regions to indicate error. The region between the inner and outer wall of the Local Bubble modelled by O’Neill et al. (2024) is indicated by the grey hatched region.
@@ 6 Discussion
In a coordinated VLBI observation of PSR B1133$+$16 we identified three arc features corresponding to distinct scintillation screens. Using both intensity and visibility data we measured the key properties of the screens: their distance, velocity, and orientation. We find that the screens are spaced roughly 50 parsecs apart. A recent survey of scintillation arcs of low dispersion measure pulsars (Stinebring et al., 2022), suggests that typical spacings between scintillation screens may be $\lesssim\!100{\rm\,pc}$, and measurement of the very nearby, bright pulsar PSR J0437--4715 reveals spacings of $\sim\!10{\rm\,pc}$ (Reardon et al., 2025). Our observations are likely consistent with these findings given observational biases (PSR B1133+16 has been observed to have at least three more screens) and sight-line dependencies (PSR B1133$+$16 is at high Galactic latitude, $69.196^{\circ}$, so may have fewer screens than pulsars closer to the Galactic plane). The screens associated with the inner and middle arc may be associated with the Local Bubble based on their distance.
@@ Acknowledgments
We gratefully acknowledge the support given by JIVE for the correlations, in particular by Bob Campbell and Benito Marcote. We also thank the Toronto scintillometry group for discussions, especially former member, Dana Simard, whose experience was very helpful in the beginning of this project. The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes. Scientific results from data presented in this publication are derived from the following EVN project code: GK049.

### v2

@@ 5.2 Association with the Local Bubble
Figure 6: The differential dust extinction as a function of distance on the line of sight towards PSR B1133+16, based on the dust map of Edenhofer et al. (2024). The distances of the screens associated with the outer arc, middle arc, and inner arc are shown with green dashdashdotted, orange dashed, and blue dashdotted lines with shaded regions to indicate error. The region between the inner and outer wall of the Local Bubble modelled by O’Neill et al. (2024) is indicated by the grey hatched region.
@@ 6 Discussion
In a coordinated VLBI observation of PSR B1133++16 we identified three arc features corresponding to distinct scintillation screens. Using both intensity and visibility data we measured the key properties of the screens: their distance, velocity, and orientation. We find that the screens are spaced roughly 50 parsecs apart. A recent survey of scintillation arcs of low dispersion measure pulsars (Stinebring et al., 2022), suggests that typical spacings between scintillation screens may be $\lesssim\!100{\rm\,pc}$, and measurement of the very nearby, bright pulsar PSR J0437-4715 reveals spacings of $\sim\!10{\rm\,pc}$ (Reardon et al., 2025). Our observations are likely consistent with these findings given observational biases (PSR B1133+16 has been observed to have at least three more screens) and sight-line dependencies (PSR B1133++16 is at high Galactic latitude, $69.196^{\circ}$, so may have fewer screens than pulsars closer to the Galactic plane).
The screens associated with the inner and middle arc may be associated with the Local Bubble based on their distance. It is unclear if or how the dust environment of a screen could manifest in other scintillation observables (e.g. persistence of screens over time or arc morphology). In this observation, the inner arc is the brightest, followed by the middle, then outer arcs. McKee et al. (2022) noted that the middle arc (their arc C) was typically the brightest and was observed most consistently. However, they suggest that all arcs persist over the 34-year data set, although they are sometimes too dim to observe. The inner arc displays arclets, but these features have also been seen in screens not associated with peaks in dust extinction.
@@ Acknowledgments
We gratefully acknowledge the support given by JIVE for the correlations, in particular by Bob Campbell and Benito Marcote. We also thank the Toronto scintillometry group for discussions, especially former member, Dana Simard, whose experience was very helpful in the beginning of this project. The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes. Scientific results from data presented in this publication are derived from the following EVN project code: GK049. We thank the anonymous referee for their comments, which helped improve the scientific quality of this work. M.H.v.K. is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) via discovery and accelerator grants.